# hard problem of consciousness

### v2

@@ 2.2 Self-Organising Systems as Self and World Constraints
Both snowflakes and human bodies are self-organising systems, yet we regard only the latter as conscious. What separates the two? To tackle this question, we need to introduce the notion of self and world constraints.
Given that we have an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the environment. Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised. We can use this fact to talk about “policies” as embodied [^7] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning. constraints on behaviour. A policy is a statement whose extension constrains outputs, like a causal intermediary in functionalism [putnam1967]. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. As a result, an organism will self-organise to express some statements, but not others. Though our environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system. This is formalised by the $\mathfrak{v}$-task as in definition 3888A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$.. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs. In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs [^9] Conversely, a system which considered all outputs to be correct would not be self-organising..
@@ 2.3 Inference
Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement. Assume we want to constrain the body to $O_{\mu}$ given $I_{\mu}$. A body could express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ could then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$. We call a constraining statement a policy. A policy constrains outputs given inputs. A correct policy is one that constrains outputs to only correct outputs. For the sake of intuition, think of “correct” as “fit” according to natural selection [^10] This need not always be the case.. This is formally expressed in definition 4.
@@
If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour. That history of interactions is just a set of statements in the organism’s embodied language, which can be expressed as a $\mathfrak{v}$-task $\mathfrak{h}$. Each input and output in $\mathfrak{h}$ is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment). The set of all fit behaviour would then be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
2.4 Learning
Learning can be understood as the process of approximating the program which caused data. Constructing a “world” model. Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$. We don’t need a world model, but a way to constrain outputs to just those that are fit. Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
@@ 2.4 Learning
Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$. These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs. Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs. Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$). If those policies imply fit behaviour given new inputs, then the organism will remain fit. If an organism learns, then it is adapting to embody a policy that it hopes will generalise to fit behaviour $\mu$.
Earlier work [bennett2023b] showed that the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions. The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [bennett2023c]. This formalised in definition 5 and illustrated by figure 1.
Figure 1: Illustration of how child tasks are examples of parent tasks.
@@ 3 Relevance Realisation Through Causal Learning
For convenience, we will now expand upon the concept of organism given in definition 6. All organisms must have preferences, because they will make some decisions and not others. Behavior implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
@@
Natural selection favours adaptability, hence we assume it prefers organisims that prefer weaker policies, generalising to remain fit. When a system optimises for weak policies, we call it weak policy optimisation (WPO). If a policy is fit, then it must correctly predict causes of valence. Hence by constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
We also need to introduce the notion of interaction between organisms. Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word). This is formally defined in 9.
@@
Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents. In that context, causality has become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust]. Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved using causal language (such as “Do Calculus” [pearl2009]) to represent an intervention from outside the system the graph describes. Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment. To illustrate this point, consider the following example from [bennett2023c].
@@
Suppose an organism named Bob is attempting to learn and predict the environment. Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now lets permit Bob to interact with its environment. Assume Bob wants it to rain. Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is absurd. $O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$. To illustrate the problem visually, we started with the acyclic graph
@@
The Psychophysical Principle of Causality
We have explained why we need to account for causality. However, we have not explained how one might come to know all the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
There are two ways to learn causal graphs. One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c]. Beginning as we are with contentless states, we do the latter. Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence. We call this the psychophysical principle of causality. This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations. We start with the complete absence of content [thompson2016], and once a system learns policies there are distinct “contents”. Policies serve to classify objects and properties. Death grounds meaning.
To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [dawid2002].
Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
@@
Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob? Couldn’t we represent the same information by introducing a new variable? The answer is yes [bennett2023c, dawid2002].
@@
Now we will show how it is possible to learn the relevant nodes to add, in the form of statements. The matter of which cause-and-effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
@@ 4 Relevant Causal Identities
The formalism we employ began as an attempt to explain why some abstractions are formed, and not others [bennettmaruyama2022a]. The optimal choice of policy for efficient adaptation is the weakest because it is implied by my statements in an embodied language. That language is dispositional, shaped by natural selection, and so a weaker policy is likely to be a fitter policy. As we have already shown, correct identification of cause and effect is often necessary to make accurate predictions and survive. WPO will thus correctly identify cause and effect where relevant. This is not just a case of substituting the “do” operator, but of simplifying the entire environment into objects and properties relevant to the organism’s motivations.
We will now explain how relevant objects are classified, as “causal identities” [bennett2023c] and protosymbols [bennett2023d].
4.1 Causal Learning
In accord with [bennett2023d, def. 6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob). A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$ in that formal language. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
@@ 4.1 Causal Learning
In other words, so long as $obs\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more interventions. For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$. If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [bennett2023c, def. 10]), giving us definition 11.
Example 1.
Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$. By affecting Alice, Bob has intervened in Alice’s experience. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
@@ 4.2 Ascribing Intent to Other Objects
The “do” operator is used to discern the difference between an event one has caused, and that same event passively observed. However, what is passive observation if not the result of an intervention by something other than one’s self? The distinction between “intervention” and not is misleading. Everything is an intervention. The question is not “is this an intervention” but “by whom was this intervention made?”.
@@
However, what if a third person Larry puts the coat on Alice? Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
@@
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. More generally, for an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity. It depends on motive, or valence. Recall that to be affected is formally defined in definition 9.
Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$. To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$). The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$. This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive. What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.

### v3

@@ 2.2 Self-Organising Systems as Self and World Constraints
Both snowflakes and human bodies self-organise, yet we regard only the latter as conscious. What separates the two? To tackle this question we need self and world constraints.
Using our embodied formal language $L_{\mathfrak{v}}$, we can talk about computation with inputs and outputs by treating everything as embodied statements embedded and enacted within the environment. Given an input $i\in L_{\mathfrak{v}}$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to those states that realise $i$, which constrains what other statements are realised. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. Hence an organism will self-organise to express some statements, but not others. Though environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system. This is formalised by the $\mathfrak{v}$-task as in definition 3777A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$.. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
@@ 2.3 Inference
Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement. Assume we want to constrain the body to $O_{\mu}$ given $I_{\mu}$. A body could express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ could then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$. We call a constraining statement a policy. A policy constrains outputs given inputs, and a correct policy is one that constrains outputs to only correct outputs. For the sake of intuition, think of “correct” as “fit” according to natural selection [^8] This need not always be the case.. This is formalised in definition 4.
2.4 Learning
If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour. That history is expressed as a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in $\mathfrak{h}$ is an interaction between organism and environment. If the set of all fit behaviour is a task $\mu$, then the organism remains fit then $\mathfrak{h}\sqsubset\mu$.
@@ 2.4 Learning
$\mathfrak{h}$ implies a set $\Pi_{\mathfrak{h}}$ of policies, and not all of those policies will constrain outputs the same way given new inputs. Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs. If those policies imply fit behaviour given new inputs, then the organism will remain fit. An organism learns fit behaviour by learning $\pi\in\Pi_{\mathfrak{h}}$ such that $\pi\in\Pi_{\mu}$.
@@
Earlier work [bennett2023b] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions. The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This means a system that does so will construct fit policies from a shorter history [bennett2023c] than one that does not (see def. 5 and proofs in appendix [bennett2024appendices]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects..
@@ 3 Relevance Realisation Through Causal Learning
All organisms (see definition 6) must have preferences, because they will make some decisions and not others. Behaviour implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
@@
Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (WPO). By constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
@@
Organisms interact when they “affect” one another (see definition 9) in the causal or physical rather than psychological sense of the word. Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents. In that context, causality has become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust]. Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [pearl2009] to represent an intervention from outside the system the graph describes.
@@
To illustrate [bennett2023c], suppose Bob is attempting to learn and predict the environment. Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now lets permit Bob to interact with its environment. Assume Bob wants it to rain. As $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is absurd. $O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$. To illustrate visually, we started with the acyclic graph
@@
3.0.1 The Psychophysical Principle of Causality
We have established the need for causal reasoning, but not how one might come to know the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
There are two ways to learn causal graphs. One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c]. We do the latter. Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence. We call this the psychophysical principle of causality. This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations. We start with the complete absence of content [thompson2016], and once a system learns policies there are distinct “contents”. Policies serve to classify objects and properties. Death grounds meaning.
To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [dawid2002]. Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
@@ 3.0.1 The Psychophysical Principle of Causality
Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From Bob’s perspective, Alice’s action assigns values to variables $A$ and $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). Likewise, we can add variables for interventions of Bob [bennett2023c].
@@
Now we will show how it is possible to learn these nodes. The matter of which cause-and-effect relations are learned is determined by valence, and so the nodes learned are statements classifying causes and valence.
@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [bennettmaruyama2022a], shaped by natural selection, and so a weaker policy is likely to be a fit policy. WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
Assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob). A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
@@
So long as $obs\neq int$, the intervention can be differentiated from the passive observation (see definition 10).
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more interventions.
For example, suppose the inputs Alice has been subject to are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $I^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $I^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-I^{\mathfrak{b}}_{\mathfrak{a}}$. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$ (see definition 11).
@@ 4.1 Ascribing Intent to Other Objects
The distinction between “intervention” and not is misleading. Passive observation is just bearing witness to an intervention by something other than one’s self. The question is not “is this an intervention" but “by whom was this intervention made?".
@@
What if a third person Larry puts the coat on Alice? Bob may observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
@@
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$. To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$). The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$. This is not to say that $\mathfrak{b}$ has intent, but intent can be ascribed to $\mathfrak{b}$ because $\mathfrak{b}$ affects $\mathfrak{a}$, who can then discern when interventions are a consequence of $\mathfrak{b}$.